Data section Manufacturing overhead cost Direct labor cost Direct labor hours Machine hours Estimated value Actual value 594,000 $ 396,000 $ 16,500 7,500 655,000 450,000 18,000 8,500 Requirement 1: Predetermined overhead rate Formula Manufacturing overhead cost/ direct labor hours Answer Applied manufacturing overhead: Formula Answer Under or over applied? Dollar amount? Requirement 2: Predetermined overhead rate Formula Answer Carry answer out to two decimal places Applied manufactring overhead Formula Answer Under or over applied? Dollar amount? Requirement 3: Predetermined overhead rate Formula Answer

Applied manufactring overhead: Formula Answer Under or over applied? Dollar amount?
Answer #1

Requirement 1:

Predetermined overhead rate(POHR)= estimated manufacturing overhead/estimated direct labour hours

POHR= 594000/16500= $36 per labour hour

Applied manufacturing overhead(AMR)= actual labour hours*POHR

AMO= 18000*36= $648000

Under applied by $7000(655000-648000)

Requirement 2:

Predetermined overhead rate= estimated manufacturing overhead/estimated machine hours

POHR= 594000/7500= $79.2 per machine hour

Applied manufacturing overhead= actual machine hours*POHR

AMO= 8500*79.2=$673200

Over applied by $18200(673200-655000)

Requirement 3:

POHR= estimated manufacturing overhead/estimated direct labour cost

= 594000/396000

=$1.5 per dollar of direct labour cost

AMO= actual direct labour cost*POHR

= 450000*1.5=$675000

Over applied by $20000 (675000-655000)

​​​​​ Note: the formula for predetermined overhead rate is (estimated manufacturing overhead/activity base). The activity base can be taken to be the estimates direct labour hours or direct labour cost or machine hours. The solution has been calculated for all the three bases.